# Formations of generalized Wannier-Stark ladders: Theorem and applications

### v1

@@ I Introduction
Rigorous results in quantum many-body systems are rare but crucial for gaining valuable insights into the characterization of dynamic behaviors within correlated systems. For non-interacting system, Wannier-Stark (WS) ladders are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field Bloch (1929); Wannier (1959, 1960); Glück (2002); Waschke et al. (1993). This has been observed in systems of semiconductor superlattices and ultracold atoms Waschke et al. (1993); Ben Dahan et al. (1996); Wilkinson et al. (1996); Anderson and Kasevich (1998). It is closely related to Bloch oscillations (BOs) due to the equidistant spectrum. In the past decade, such a phenomenon has attracted much attention in cold-atoms physics and photonics due to applications in interferometric measurements and as a method for manipulating localized wave packets Breid et al. (2006, 2007); Dreisow et al. (2009); Kling et al. (2010); Plötz and Wimberger (2011). It can be simulated using artificial quantum systems, such as superconducting circuits Song et al. (2024). On the other hand, the dynamics of particle pairs in lattice systems have garnered considerable interest, owing to the rapid advancements in experimental techniques. Ultra-cold atoms have proven to be an ideal testing ground for few-particle fundamental physics, as optical lattices offer clean realizations of a variety of many-body Hamiltonians. It stimulates many experimental Winkler et al. (2006); Fölling et al. (2007); Gustavsson et al. (2008) and theoretical investigations Mahajan and Thyagaraja (2006); Petrosyan et al. (2007); Creffield (2007); Kuklov and Moritz (2007); Zöllner et al. (2008); Wang et al. (2008); Valiente and Petrosyan (2008); Jin et al. (2009); Valiente and Petrosyan (2009); Valiente et al. (2010); Javanainen et al. (2010); Wang and Liang (2010); Rosch et al. (2008); Zhang (2024) in strongly correlated systems. Recently, it has been shown that a bound pair within a correlated system also exhibits periodic dynamics when subjected to a linear potential Khomeriki et al. (2010); Longhi (2011, 2012); Corrielli et al. (2013); Lin et al. (2014); Zhang et al. (2016); Zhang (2024).
@@
In this work, we extend the concept of WSL to the many-body systems. We investigate the formations of WSL in generalized systems, including strongly correlated and non-Hermitian systems. We present a theorem on the existence of WSL for a set of general systems that are translationally symmetric before the addition of a linear potential. We show that the WSL becomes complex but maintains a real energy level spacing for a non-Hermitian system. To illustrate the theorem, we focus on 1D extended Bose-Hubbard models with both real and imaginary hopping strengths. It is shown that the Bloch-Zener oscillations of correlated bosons are particularly remarkable under resonant conditions. Numerical simulations for cases with boson numbers $n=2$, $3$, and $4$ are presented. Analytical and numerical results for the time evolution of the $n$-boson-occupied initial state indicate that all evolved states exhibit quasi periodic oscillations, but with different profiles, depending on the Hermiticity and interaction strength.
@@
This paper is organized as follows. In Sec. II, we present the theorem on the structure of energy levels for general systems in a linear field, including Hermitian and non-Hermitian. In Sec. III, we take the extended Bose-Hubbard model as an application of the theorem and provide its effective Hamiltonian. Sec. IV and Sec. V are dedicated to the dynamics of the multi-particle effective Hamiltonian and original Hamiltonian, respectively. Finally, we provide a summary in Sec. VI.
@@ II Theorem on WSLs
$\left[T_{n_{0}},H_{0}\right]=0,$ (3)
where $T_{n_{0}}$ is the translational operator defined as
$T_{n_{0}}a_{j,\sigma}T_{n_{0}}^{-1}=a_{j+n_{0},\sigma}.$ (4)
@@
$\left[T_{r_{0}},H\right]=-r_{0}\omega\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma},$ (5)
which is referred to as a ramped translational symmetry. in the invariant subspace with fixed particle number $n_{0}$, any eigenstate of $H$ obeys
@@
$\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}|\psi_{0}\rangle=n_{0}|\psi_{0}\rangle.$ (7)
@@
$H\left(T_{r_{0}}|\psi_{0}\rangle\right)=\left(E_{0}+r_{0}\mathcal{\omega}\right)\left(T_{r_{0}}|\psi_{0}\rangle\right),$ (8)
@@
$H\left(T_{r_{0}}^{-1}|\psi_{0}\rangle\right)=\left(E_{0}-r_{0}\mathcal{\omega}\right)\left(T_{r_{0}}|\psi_{0}\rangle\right),$ (9)
i.e., $T_{r_{0}}|\psi_{0}\rangle$ ($T_{r_{0}}^{-1}|\psi_{0}\rangle$) is also the eigenstate of $H$ with eigen energy $E_{0}+r_{0}\mathcal{\omega}$ ($E_{0}-r_{0}\mathcal{\omega}$). Operator $T_{r_{0}}$ increments the energy level by a step of $r_{0}\mathcal{\omega}$ and the operator $T_{r_{0}}^{-1}$ decrements the energy level by a step of $r_{0}\mathcal{\omega}$. We can then construct a set of eigenstates
$|\psi_{m}\rangle=\left(T_{r_{0}}\right)^{m}|\psi_{0}\rangle,$ (10)
@@
$E_{m}=E_{0}+mr_{0}\mathcal{\omega}.$ (11)
It indicates that the translational operator $T_{r_{0}}$ acts as a ladder operator. This proof is independent of $H_{0}$. We note that based on another eigenstate $|\psi_{0}^{\prime}\rangle$, which does not be long to $\left\{|\psi_{m}\rangle\right\}$, another set of eigenstates $\left\{|\psi_{m}^{\prime}\rangle\right\}$ can be generated accordingly. Although $E_{0}$ can be real or complex, dependenting on the detail of the Hamiltonian $H_{0}$, we can conclude that the energy levels of Hamiltonian $H$ consists of multi-set of WSLs with an identical real level spacing $r_{0}\mathcal{\omega}$.
@@
This rigorous conclusion has important implications for investigating the dynamics of many-body and non-Hermitian systems in the presence of a linear field. In the following sections, we will focus on both the Hermitian and non-Hermitian extended Bose-Hubbard models to demonstrate the application of our conclusions and to reveal the dynamics of correlated bosons.
@@ III Extended Bose-Hubbard model
where $b_{j}^{{\dagger}}$ and $n_{j}=b_{j}^{{\dagger}}b_{j}$ is the creation and number operators of the boson at the $j$th site. Parameters $\kappa$, $U$ and $V$, denote the tunneling strength, on-site and NN interactions between bosons. In this work, we consider the cases with real and imaginary , respectively. Obviously, Hamiltonian $H_{0}$ satisfies the conditions in the theorem, and then possesses multi-set of WSLs.
@@ III.2 Resonant interactions
$\left|jn+\lambda\right\rangle=\frac{\left(b_{j}^{\dagger}\right)^{n-\lambda}\left(b_{j+1}^{\dagger}\right)^{\lambda}}{\sqrt{\left(n+1-\lambda\right)!\left(\lambda-1\right)!}}|\mathrm{vac}\rangle,$ (16)
@@
In order to obtain the effective Hamiltonian, one can simply envalue the nonzero matrix elements of $H_{\text{{B}}}$ in the subspace spanned by the set of eigenstates $\left\{\left|l\right\rangle\right\}$. For a given boson number $n$, the effective Hamiltonian can be written as
@@
In its present form, $H_{\mathrm{eff}}^{[n]}$ are formally analogous to a tight-binding model describing a single-particle dynamics in an infinite chain with NN hopping, under a linear field. We note that the states $\left|l\right\rangle$ in each unit cell represent an $n$-bosonic-cluster state, in which all $n$ bosons occupy a single dimer of the original lattice of the Bose-Hubbard system $H_{\text{{B}}}$. In this sense, the dynamics of $H_{\mathrm{eff}}^{[n]}$ is essentially the dynamics of $n$ correlated bosons. Notably, we would like to point out that the hopping strength in $H_{\mathrm{eff}}^{[n]}$ is of the order of $\kappa$, not $\kappa^{2}/U$. This ensures that the phenomena arising from the effective Hamiltonian can be observed in experiments, similar to those of a single boson. On the other hand, our theorem also holds for the effective Hamiltonian $H_{\mathrm{eff}}^{[n]}$, which describes a single-particle system of $n$-site unit cell. The corresponding step for each ladder is $n\omega$.
@@ IV Dynamics of effective Hamiltonian $H_{\mathrm{eff}}^{[n]}$
$+\sqrt{6}|4j+2\rangle\langle 4j+3|\,+2|4j+3\rangle\langle 4j+4|+\mathrm{H.c.})$
Numerical simulation shows that, when $\kappa$ is imaginary, there are two sets of real ladders and two sets of complex ladders. The two real ladders form a new ladder with energy levels given by ±$\pm 2\omega m$, where $m=0$, $1$, …. The real parts of two complex ladders is given by $(2\pm 4m)\omega$, where $m=0$, $1$, ….
@@ V Dynamics of correlated bosons
In this section, we focus on the investigation on the dynamics of correlated bosons. Considering the original Bose-Hubbard Hamiltonian as given by Eq. (12), the theorem regarding its energy levels applies to any set of parameters $\kappa$, $U$ and $V$. This feature can be characterized by the quantity $\varepsilon$, which is defined as the sum of the real part of $E$ and the imaginary part of $E$,
@@ VI Summary
In summary, we have developed a theory for a class of Hermitian and non-Hermitian Hamiltonians that possess WSLs. Such Hamiltonians consist of two distinct parts: an arbitrary Hamiltonian exhibiting translational symmetry and a linear potential. The most fascinating and important feature of these systems is that their conclusions are independent of the details of the Hamiltonian. As examples, we have investigate the extended Bose-Hubbard model with several bosons. Analytic and numerical approaches verified our conclusion in the case with $U=V$. We also demonstrated the application of the theorem by analyzing dynamics driven by the Hubbard model using numerical simulations. Our findings provide a method for analyzing the energy-level structure of general systems and are expected to be both necessary and insightful for quantum engineering in many-body systems.

### v2

@@ I Introduction
Rigorous results in quantum interacting systems are rare but crucial for gaining valuable insights into the characterization of dynamic behaviors within correlated systems. For non-interacting system, Wannier-Stark ladders (WSL) are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field Bloch (1929); Wannier (1959, 1960); Glück (2002); Waschke et al. (1993). This has been observed in systems of semiconductor superlattices and ultracold atoms Waschke et al. (1993); Ben Dahan et al. (1996); Wilkinson et al. (1996); Anderson and Kasevich (1998). It is closely related to Bloch oscillations (BOs) due to the equidistant spectrum. In the past decade, such a phenomenon has attracted much attention in cold-atoms physics and photonics due to applications in interferometric measurements and as a method for manipulating localized wave packets Breid et al. (2006, 2007); Dreisow et al. (2009); Kling et al. (2010); Plötz and Wimberger (2011). It can be simulated using artificial quantum systems, such as superconducting circuits Song et al. (2024). On the other hand, the dynamics of particle pairs in lattice systems have garnered considerable interest, owing to the rapid advancements in experimental techniques. Ultra-cold atoms have proven to be an ideal testing ground for few-particle fundamental physics, as optical lattices offer clean realizations of a variety of interacting Hamiltonian systems. It stimulates many experimental Winkler et al. (2006); Fölling et al. (2007); Gustavsson et al. (2008) and theoretical investigations Mahajan and Thyagaraja (2006); Petrosyan et al. (2007); Creffield (2007); Kuklov and Moritz (2007); Zöllner et al. (2008); Wang et al. (2008); Valiente and Petrosyan (2008); Jin et al. (2009); Valiente and Petrosyan (2009); Valiente et al. (2010); Javanainen et al. (2010); Wang and Liang (2010); Rosch et al. (2008); Zhang (2024) in strongly correlated systems. Recently, it has been shown that a bound pair within a correlated system also exhibits periodic dynamics when subjected to a linear potential Khomeriki et al. (2010); Longhi (2011, 2012); Corrielli et al. (2013); Lin et al. (2014); Zhang et al. (2016); Zhang (2024).
@@
In this work, we extend the concept of WSL to the interacting Hamiltonian systems. We investigate the formations of WSL in generalized systems, including strongly correlated and non-Hermitian systems. We present a theorem on the existence of WSL for a set of general systems that are translationally symmetric before the addition of a linear potential. We show that the WSL becomes complex but maintains a real energy level spacing for a non-Hermitian system. To illustrate the theorem, we focus on 1D extended Bose-Hubbard models with both real and imaginary hopping strengths. It is shown that the Bloch-Zener oscillations of correlated bosons are particularly remarkable under resonant conditions. Numerical simulations for cases with boson numbers $N=2$, $3$, and $4$ are presented. Analytical and numerical results for the time evolution of the $N$-boson-occupied initial state indicate that all evolved states exhibit quasi periodic oscillations, but with different profiles, depending on the Hermiticity and interaction strength.
@@
This paper is organized as follows. In Sec. II, we present the theorem on the structure of energy levels for general systems in a linear field, including Hermitian and non-Hermitian. In Sec. III, we take the extended Bose-Hubbard model as an application of the theorem and provide its effective Hamiltonian under the resonant condition in strongly interaction limit. Sec. IV dedicated to the dynamics of the few-boson effective Hamiltonian and the corresponding original Hamiltonian, respectively. In particular, we also provide numerical simulations for both Hermitian and non-Hermitian sample systems, which offer evident demonstrations of our rigorous results, in Sec. V. Finally, we provide a summary in Sec. VI.
@@ II Theorem on WSLs
$\left[T_{r},H_{0}\right]=0,$ (3)
where $T_{r}$ is the translational operator defined as
$T_{r}a_{j,\sigma}T_{r}^{-1}=a_{j+r,\sigma}.$ (4)
@@
$\left[T_{r},H\right]=-r\omega\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}T_{r},$ (5)
which is referred to as a ramped translational symmetry. in the invariant subspace with fixed particle number $N$, any eigenstate of $H$ obeys
@@
$\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}|\psi_{0}\rangle=N|\psi_{0}\rangle.$ (7)
@@
$H\left(T_{r}|\psi_{0}\rangle\right)=\left(E_{0}+Nr\mathcal{\omega}\right)\left(T_{r}|\psi_{0}\rangle\right),$ (8)
@@
$H\left(T_{r}^{-1}|\psi_{0}\rangle\right)=\left(E_{0}-Nr\mathcal{\omega}\right)\left(T_{r}^{-1}|\psi_{0}\rangle\right),$ (9)
i.e., $T_{r}|\psi_{0}\rangle$ ($T_{r}^{-1}|\psi_{0}\rangle$) is also the eigenstate of $H$ with eigen energy $E_{0}+Nr\mathcal{\omega}$ ($E_{0}-Nr\mathcal{\omega}$). Operator $T_{r}$ increments the energy level by a step of $Nr\mathcal{\omega}$ and the operator $T_{r}^{-1}$ decrements the energy level by a step of $Nr\mathcal{\omega}$. We can then construct a set of eigenstates
$|\psi_{m}\rangle=\left(T_{r}\right)^{m}|\psi_{0}\rangle,$ (10)
@@
$E_{m}=E_{0}+mNr\mathcal{\omega}.$ (11)
It indicates that the translational operator $T_{r}$ acts as a ladder operator. This proof is independent of $H_{0}$. We note that based on another eigenstate $|\psi_{0}^{\prime}\rangle$, which does not be long to $\left\{|\psi_{m}\rangle\right\}$, another set of eigenstates $\left\{|\psi_{m}^{\prime}\rangle\right\}$ can be generated accordingly. Although $E_{0}$ can be real or complex, dependenting on the detail of the Hamiltonian $H_{0}$, we can conclude that the energy levels of Hamiltonian $H$ consists of multi-set of WSLs with an identical real level spacing $Nr\mathcal{\omega}$. In principle, this conclusion is true for any given $N$. However, a large energy level spacing that exceeds the scale of low-energy physics is meaningless.
Fortunately, this conclusion drawn from few-body systems is applicable to specific invariant subspaces of many-body systems, in which energy eigenstates are localized. Here, we provide an example to demonstrate this point. Consider a system with a set of $n$-boson cluster eigenstates in the $n$-boson invariant subspace. These eigenstates are connected by the translational operator. In this situation, one could easily have a large number $K$ of widely separated $n$-boson clusters to construct the eigenstates with a total boson number $N=Kn$. The WSL frequency spacing derived above for states of that multi-cluster kind would therefore be $Knr\omega$. However, there exist many subsets of eigenstates, which can be constructed by translating one $n$-boson cluster while leaving the other $K-1$ clusters unchanged. Then, the corresponding energy levels in a subset form the WSL with frequency spacing $nr\omega$ rather than $Nr\omega$. This indicates that a system with a large boson number $N=Kn$ can also support dynamics with a smaller frequency $nr\omega$.
@@
This rigorous conclusion has important implications for investigating the dynamics of interacting and non-Hermitian systems in the presence of a linear field. In the following sections, we will focus on both the Hermitian and non-Hermitian extended Bose-Hubbard models to demonstrate the application of our conclusions and to reveal the dynamics of correlated bosons.
@@ III Extended Bose-Hubbard model
where $b_{j}^{{\dagger}}$ and $n_{j}=b_{j}^{{\dagger}}b_{j}$ is the creation and number operators of the boson at the $j$th site. Parameters $\kappa$, $U$ and $V$, denote the tunneling strength, on-site and NN interactions between bosons. In this work, we consider the cases with real and imaginary $\kappa$, respectively. Obviously, Hamiltonian $H_{0}$ satisfies the conditions in the theorem, and then possesses multi-set of WSLs.
@@ III.2 Resonant interactions
$\left|jN+\lambda\right\rangle=\frac{\left(b_{j}^{\dagger}\right)^{N-\lambda}\left(b_{j+1}^{\dagger}\right)^{\lambda}}{\sqrt{\left(N-\lambda\right)!\left(\lambda\right)!}}|\mathrm{vac}\rangle,$ (16)
@@
In order to obtain the effective Hamiltonian, one can simply evaluate the nonzero matrix elements of $H_{\text{{B}}}$ in the subspace spanned by the set of eigenstates $\left\{\left|l\right\rangle\right\}$. For a given boson number $N$, the effective Hamiltonian can be written as
@@
In its present form, $H_{\mathrm{eff}}^{[N]}$ are formally analogous to a tight-binding model describing a single-particle dynamics in an infinite chain with NN hopping, under a linear field. We note that the states $\left|l\right\rangle$ in each unit cell represent an $N$-bosonic-cluster state, in which all $N$ bosons occupy a single dimer of the original lattice of the Bose-Hubbard system $H_{\text{{B}}}$. In this sense, the dynamics of $H_{\mathrm{eff}}^{[N]}$ is essentially the dynamics of $N$ correlated bosons. In Fig. 1, a schematic illustration showing the correspondence between the configurations of the boson cluster with $N=3$ and the basis states for the effective Hamiltonian under strong interactions is presented.
Notably, we would like to point out that the hopping strength in $H_{\mathrm{eff}}^{[N]}$ is of the order of $\kappa$, not $\kappa^{2}/U$. This ensures that the phenomena arising from the effective Hamiltonian can be observed in experiments, similar to those of a single boson. On the other hand, our theorem also holds for the effective Hamiltonian $H_{\mathrm{eff}}^{[N]}$, which describes a single-particle system of $N$-site unit cell. The corresponding step for each ladder is $N\omega$.
Figure 1: Schematic illustrations of the correspondence between the configurations of the boson cluster with $N=3$ and the basis states for the effective Hamiltonian given in Eq. 17. (a1-a4) Four configurations of the boson cluster with $N=3$. These are degenerate states of the Hamiltonian $H_{0}(\kappa=0)$. Here, only the hopping processes within a dimer are presented. (b1-b4) The sites in the chain represent the basis states of the effective Hamiltonian. The hopping strengths indicated are non-uniform.
@@ IV Dynamics of effective Hamiltonian $H_{\mathrm{eff}}^{[N]}$
$+\sqrt{6}|4j+2\rangle\langle 4j+3|\,+2|4j+3\rangle\langle 4j+4|$
$+\mathrm{H.c.})+\omega\sum_{l=-\infty}^{\infty}l\left|l\right\rangle\left\langle l\right|.$
Numerical simulation shows that, when $\kappa$ is imaginary, there are two sets of real ladders and two sets of complex ladders. The two real ladders form a new ladder with energy levels given by ±$\pm 2m\omega$, where $m=0$, $1$, …. The real parts of two complex ladders is given by $(2\pm 4m)\omega$, where $m=0$, $1$, ….
@@ V Dynamics of correlated bosons
According to the theorem, WSLs exist in a Bose-Hubbard model with arbitrary parameters $\kappa$, $U$ and $V$. However, the numerical results presented so far have been essentially for single-particle systems. In this section, we will focus on both the Hermitian and non-Hermitian extended Bose-Hubbard models with moderate values of $U$ and $V$. We will provide numerical results to show the existence of WSLs and present numerical results on the dynamics of correlated bosons, which offer evident demonstrations of our rigorous results.
The feature of WSLs can be characterized by the quantity $\varepsilon$, which is defined as the sum of the real part of $E$ and the imaginary part of $E$,
@@ VI Summary
In summary, we have developed a theory for a class of Hermitian and non-Hermitian Hamiltonians that possess WSLs. Such Hamiltonians consist of two distinct parts: an arbitrary Hamiltonian exhibiting translational symmetry and a linear potential. The results are significant for few-body systems or for many-body systems with few-body clusters. As examples, we have investigate the extended Bose-Hubbard model with several bosons. Analytic and numerical approaches verified our conclusion in the cases with large and moderate $U=V$, real and imaginary $\kappa$. We also demonstrated the application of the theorem by analyzing dynamics driven by the Hubbard model using numerical simulations. Our findings provide a method for analyzing the energy-level structure of general systems and are expected to be both necessary and insightful for quantum engineering in quantum interacting systems.